# Kengo Nakamura, Masaaki Nishino, and Shuhei Denzumi NTT Communication Science Laboratories, Kyoto, Japan

### v1

@@ 2.3 Family Algebra Operations on ZDDs
In this section, we explain how we the family algebra operations are performed using ZDDs and point out what makes the difference between the basic set operations (union, intersection, and difference) and the other operations.
@@
The case of $\mathsf{lb}(\mathtt{m})<\mathsf{lb}(\mathtt{n})$ can be handled in the same way. By recursively expanding $\mathcal{F}_{\mathtt{n}}\cup\mathcal{F}_{\mathtt{m}}$ by (1) and (2), we eventually reach terminal nodes where the union is trivial, e.g., $\mathcal{F}_{\bot}\cup\mathcal{F}_{\top}=\{\emptyset\}$. Therefore, by caching the resultant ZDD nodes of $\mathcal{F}_{\mathtt{n}^{\prime}}\cup\mathcal{F}_{\mathtt{m}^{\prime}}$, where $\mathtt{n}^{\prime}$ and $\mathtt{m}^{\prime}$ are the child nodes of $\mathtt{n}$ and $\mathtt{m}$, respectively, we can efficiently compute the ZDD representing $\mathcal{F}_{\mathtt{n}}\cup\mathcal{F}_{\mathtt{m}}$. With the cache, one can show that we can build a ZDD representing the union of two ZDDs in a time proportional to the product of input ZDD sizes. The intersection and difference operations can be handled in almost the same way.
The other operations can also be performed in a recursive manner. However, the recursion procedure is more complicated than that of the basic set operations. Let us consider, for example, the join operation. Given $\mathcal{F}_{\mathtt{n}}$ and $\mathcal{F}_{\mathtt{m}}$, where $\mathsf{lb}(\mathtt{n})=\mathsf{lb}(\mathtt{m})=x$, their join can be represented as follows:
@@
$=$ $[\mathcal{F}_{\mathsf{lo}(\mathtt{n})}\sqcup\mathcal{F}_{\mathsf{lo}(\mathtt{m})}]\cup$
[$\ [\{\{x\}\}\sqcup((\mathcal{F}_{\mathsf{lo}(\mathtt{n})}\sqcup\mathcal{F}_{\mathsf{hi}(\mathtt{m})})\cup(\mathcal{F}_{\mathsf{hi}(\mathtt{n})}\sqcup\mathcal{F}_{\mathsf{lo}(\mathtt{m})})\cup(\mathcal{F}_{\mathsf{hi}(\mathtt{n})}\sqcup\mathcal{F}_{\mathsf{hi}(\mathtt{m})}))].$
@@
$\mathcal{F}_{\mathtt{n}}\bigtriangleup\mathcal{F}_{\mathtt{m}}=$ $[\mathcal{F}_{\mathsf{lo}(\mathtt{n})}\bigtriangleup\mathcal{F}_{\mathsf{lo}(\mathtt{m})}]\cup$ (4)
[$\ [\{\{x\}\}\sqcup(\mathcal{F}_{\mathsf{hi}(\mathtt{n})}\bigtriangleup(\mathcal{F}_{\mathsf{lo}(\mathtt{m})}\cup\mathcal{F}_{\mathsf{hi}(\mathtt{m})}))].$
@@ 3.1 High-Level Idea
We use such families to constitute examples of blow-up. More specifically, for each operation, we constitute an example such that performing this operation incurs the union or intersection of multiple families. Since we prove that the reduced ZDD representing the result of an operation will become exponential in size, we can confirm that any algorithm for computing the resultant ZDD incurs worst-case non-polynomial complexity. Combined with concrete instances where multiple union or intersection operations lead to exponential blow-up, we prove that the worst-case complexity of family algebra operations is lower-bounded by an exponential factor.
@@
The hidden weighted bit function $\mathcal{H}_{m}$ can be represented as a union of elementary families. Define $\mathcal{E}_{m,k}\coloneqq\{S\subseteq\{y_{1},\ldots,y_{m}\}\mid|S|=k,y_{k}\in S\}$, i.e., $\mathcal{E}_{m,k}$ consists of the subsets of $\{y_{1},\ldots,y_{m}\}$ where the cardinality is $k$ and $y_{k}$ is contained. Then, $\mathcal{H}_{m}=\bigcup_{k=1}^{m}\mathcal{E}_{m,k}$. It can be easily verified that the size of the ZDD representing $Z(\mathcal{E}_{m,k})=O(m^{2})$ for any order of elements (see Appendix). However, it is proved that the ZDD representing $\mathcal{H}_{m}$ must become exponential in size.
Theorem 3 ([2]).
For any order $<$ of elements, $Z_{<}(\mathcal{H}_{m})=\Omega(2^{m/5}/m)$.
Definition 4.
@@
For $k=1,\ldots,m$, let $\mathcal{Q}_{m,k}$ be the family of subsets of $\{y_{1},\ldots,y_{m^{2}}\}$ such that there is exactly one element from $y_{m(k\!-\!1)\!+\!1},y_{m(k\!-\!1)\!+\!2},\ldots,y_{m(k\!-\!1)\!+\!m}$, and let $\mathcal{Q}_{m,m+k}$ be those such that there is exactly one element from $y_{k},y_{m\!+\!k},\ldots,y_{m(m\!-\!1)\!+\!k}$. Then, $\mathcal{P}_{m}=\bigcap_{k=1}^{2m}\mathcal{Q}_{m,k}$. Here, $Z(\mathcal{Q}_{m,k})=O(m^{2})$ for any order of elements, as proved in Appendix. However, it is again proved that the ZDD representing $\mathcal{P}_{m}$ must become exponential in size.
Theorem 5 ([6]).
For any order $<$ of elements, $Z_{<}(\mathcal{P}_{m})=\Omega(2^{m}/m)$.
We first show the exponential blow-up cases for a specific order of elements in Section 3.2. However, we see that the size of ZDD representing the hidden weighted bit function or the permutation function is exponential regardless of the order of elements. Therefore, in Section 3.3, we prove that for each family generated by the operation in Section 3.2, the ZDD size representing it remains exponential regardless of the order of elements. This means that for each operation, there exists an instance in which the input ZDD size can be polynomial by manipulating the element order but the output ZDD size must be exponential for any element order.
@@ 3.2 Proofs with Specific Element Order
The theorems in this subsection are proved under a specific order of elements for each.
@@ 3.2.1 Join, Disjoint Join, Joint Join, Meet, and Delta
We first treat the following five operations: join, disjoint join, joint join, meet, and delta. For these operations, we constitute a pair of families such that performing an operation incurs the union of $O(m)$ subfamilies. Combined with $\mathcal{E}_{m,k}$, the result after taking an operation contains $\bigcup_{k}\mathcal{E}_{m,k}=\mathcal{H}_{m}$, which is the hidden weighted bit function for which the ZDD size is exponential in $m$.
Theorem 6.
@@
Figure 2: Example of blow-up for join operation. Blue triangles mean that the ZDD size representing this family is polynomial in $m$, while red triangle means that its size is exponential in $m$. Arcs going to $\bot$ terminal are omitted.
Proof.
Let us consider the families of subsets of $X\cup Y$, where $X\coloneqq\{x_{1},\ldots,x_{m}\}$ and $Y\coloneqq\{y_{1},\ldots,y_{m}\}$. The order of elements is determined as $x_{1},\ldots,x_{m},y_{1},\ldots,y_{m}$. We define $\mathcal{F}_{m}$ as
@@
$\mathcal{F}_{m}\sqcup\mathcal{G}_{m}$ $\textstyle=(\bigcup_{k=1}^{m}(\{\{x_{k}\}\}\sqcup\mathcal{E}_{m,k}))\sqcup\{X\}$
$\textstyle=\bigcup_{k=1}^{m}((\{\{x_{k}\}\}\sqcup\mathcal{E}_{m,k}))\sqcup\{X\})$
$\textstyle=\bigcup_{k=1}^{m}(\{X\}\sqcup\mathcal{E}_{m,k}))$
=$\textstyle=\{X\}\sqcup(\bigcup_{k=1}^{m}\mathcal{E}_{m,k})=\{X\}\sqcup\mathcal{H}_{m},$
@@
$\mathcal{F}_{m}\operatorname{\dot{\bowtie}}\mathcal{G}_{m}$ $\textstyle=\bigcup_{k=1}^{m}((\{x_{k}\}\cup(X\setminus\{x_{k}\}))\sqcup\mathcal{E}_{m,k})$
=$\textstyle=\{X\}\sqcup(\bigcup_{k=1}^{m}\mathcal{E}_{m,k})=\{X\}\sqcup\mathcal{H}_{m},$
@@
For the delta operation, we let $\mathcal{G}_{m}=2^{X}$. Since $\{x_{k}\}\operatorname{\boxplus}2^{X}=2^{X}$ for any $k$, we have
$\mathcal{F}_{m}\operatorname{\boxplus}\mathcal{G}_{m}$ $\textstyle=\bigcup_{k=1}^{m}((\{x_{k}\}\operatorname{\boxplus}2^{X})\sqcup\mathcal{E}_{m,k})$
$\textstyle=2^{X}\sqcup(\bigcup_{k=1}^{m}\mathcal{E}_{m,k})=2^{X}\sqcup\mathcal{H}_{m}.$
The ZDD size of $\mathcal{F}_{m}\operatorname{\boxplus}\mathcal{G}_{m}$ is at least $Z(\mathcal{H}_{m})=\Omega(2^{m/5}/m)$. ∎
@@ 3.2.2 Quotient and Remainder
Next, we focus on the quotient and remainder operations. For the quotient operation, we constitute a pair of families such that performing an operation incurs the intersection of $O(m)$ subfamilies. Here, let $\mathcal{E}^{\prime}_{m,k}\coloneqq 2^{Y}\setminus\mathcal{E}_{m,k}$ be the complement of $\mathcal{E}_{m,k}$ regarding the family of subsets of $Y$. By De Morgan’s laws, we have $\bigcap_{k}\mathcal{E}^{\prime}_{m,k}=2^{Y}\setminus(\bigcup_{k}\mathcal{E}_{m,k})=2^{Y}\setminus\mathcal{H}_{m}\eqqcolon\mathcal{H}^{\prime}_{m}$. The ZDD size representing $\mathcal{H}^{\prime}_{m}$ can be lower bounded by the following lemma.
Lemma 7.
@@
Proof.
$\mathcal{F}\supseteq\mathcal{G}$ implies $\mathcal{F}\setminus(\mathcal{F}\setminus\mathcal{G})=\mathcal{G}$. Since the ZDD size after taking the difference can be bounded by the product of the sizes of operand ZDDs, we have $Z(\mathcal{G})=O(Z(\mathcal{F})Z(\mathcal{F}\setminus\mathcal{G}))$. Suppose $Z(\mathcal{F}\setminus\mathcal{G})=o(g(m)/f(m))$. Then, $Z(\mathcal{G})=o(f(m)\cdot(g(m)/f(m)))=o(g(m))$, refuting the assumption $Z(\mathcal{G})=\Omega(g(m))$. Therefore, $Z(\mathcal{F}\setminus\mathcal{G})=\Omega(g(m)/f(m))$. ∎
Since $Z(2^{Y})=O(m)$ and $Z(\mathcal{H}_{m})=\Omega(2^{m/5}/m)$, we have $Z(\mathcal{H}^{\prime}_{m})=\Omega(2^{m/5}/m^{2})$. The remainder operation can also be lower bounded using Lemma 7.
Theorem 8.
@@
Figure 3: Example of blow-up for quotient operation.
Proof.
@@
We have $Z(\mathcal{E}^{\prime}_{m,k})=O(m^{2})$ as proved in Appendix, and thus $Z(\mathcal{F}_{m})=O(m^{3})$. We also define $\mathcal{G}_{m}\coloneqq\{\{x_{1}\},\ldots,\{x_{m}\}\}$, where $Z(\mathcal{G}_{m})=O(m)$.
@@
For the remainder operation, we prepared the same families. Since $\mathcal{G}_{m}\sqcup(\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m})=\{\{x_{1}\},\ldots,\{x_{m}\}\}\sqcup\mathcal{H}^{\prime}_{m}$, $Z(\mathcal{G}_{m}\sqcup(\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m}))=\Omega(2^{m/5}/m^{2})$. Also, since $S\in\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m}$ if and only if $S\cup G\in\mathcal{F}_{m}$ for all $G\in\mathcal{G}_{m}$, all of the subsets in $\mathcal{G}_{m}\sqcup(\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m})$ are also contained in $\mathcal{F}_{m}$. In other words, $\mathcal{F}_{m}\supseteq\mathcal{G}_{m}\sqcup(\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m})$. Therefore, by using Lemma 7, $Z(\mathcal{F}_{m}\operatorname{\%}\mathcal{G}_{m})=\Omega((2^{m/5}/m^{2})/m^{3})=\Omega(2^{m/5}/m^{5})$. ∎

### v2

@@ 2.2 Family Algebra Operations on ZDDs
In this section, we explain how the family algebra operations are performed using ZDDs and point out what makes the difference between the basic set operations (union, intersection, difference, and symmetric difference) and the other operations.
@@
The case of $\mathsf{lb}(\mathtt{m})<\mathsf{lb}(\mathtt{n})$ can be handled in the same way. By recursively expanding $\mathcal{F}_{\mathtt{n}}\cup\mathcal{F}_{\mathtt{m}}$ by (1) and (2), we eventually reach terminal nodes where the union is trivial, e.g., $\mathcal{F}_{\bot}\cup\mathcal{F}_{\top}=\{\emptyset\}$. Therefore, by caching the resultant ZDD nodes of $\mathcal{F}_{\mathtt{n}^{\prime}}\cup\mathcal{F}_{\mathtt{m}^{\prime}}$, where $\mathtt{n}^{\prime}$ and $\mathtt{m}^{\prime}$ are the child nodes of $\mathtt{n}$ and $\mathtt{m}$, respectively, we can efficiently compute the ZDD representing $\mathcal{F}_{\mathtt{n}}\cup\mathcal{F}_{\mathtt{m}}$. With the cache, one can show that we can build a ZDD representing the union of two ZDDs in a time proportional to the product of input ZDD sizes. The intersection, difference, and symmetric difference operations can be handled in almost the same way.
The other operations can also be performed in a recursive manner. However, the recursion becomes more complicated. Let us consider, for example, the join operation. When $\mathsf{lb}(\mathtt{n})=\mathsf{lb}(\mathtt{m})=x$, the join becomes
@@
$=$ $[\mathcal{F}_{\mathsf{lo}(\mathtt{n})}\!\sqcup\!\mathcal{F}_{\mathsf{lo}(\mathtt{m})}]\!\cup\![\{\{x\}\}\sqcup((\mathcal{F}_{\mathsf{lo}(\mathtt{n})}\!\sqcup\!\mathcal{F}_{\mathsf{hi}(\mathtt{m})})\cup(\mathcal{F}_{\mathsf{hi}(\mathtt{n})}\!\sqcup\!\mathcal{F}_{\mathsf{lo}(\mathtt{m})})\cup(\mathcal{F}_{\mathsf{hi}(\mathtt{n})}\!\sqcup\!\mathcal{F}_{\mathsf{hi}(\mathtt{m})}))].$
@@
$\mathcal{F}_{\mathtt{n}}\bigtriangleup\mathcal{F}_{\mathtt{m}}=$ $[\mathcal{F}_{\mathsf{lo}(\mathtt{n})}\bigtriangleup\mathcal{F}_{\mathsf{lo}(\mathtt{m})}]\cup[\{\{x\}\}\sqcup(\mathcal{F}_{\mathsf{hi}(\mathtt{n})}\bigtriangleup(\mathcal{F}_{\mathsf{lo}(\mathtt{m})}\cup\mathcal{F}_{\mathsf{hi}(\mathtt{m})}))].$ (4)
@@ 3.1 High-Level Idea
We use such families to constitute examples of blow-up. More specifically, for each operation, we constitute an example such that performing this operation incurs the union or intersection of multiple families. Since we prove that the reduced ZDD representing the result of an operation will become exponential in size, we can confirm that any algorithm for computing the resultant ZDD incurs worst-case non-polynomial complexity. Combined with concrete instances, we prove that the worst-case complexity of family algebra operations is lower-bounded by an exponential factor.
@@
The hidden weighted bit function $\mathcal{H}_{m}$ can be represented as a union of elementary families. Define $\mathcal{E}_{m,k}\coloneqq\{S\subseteq\{y_{1},\ldots,y_{m}\}\mid|S|=k,y_{k}\in S\}$, i.e., $\mathcal{E}_{m,k}$ consists of the subsets of $\{y_{1},\ldots,y_{m}\}$ where the cardinality is $k$ and $y_{k}$ is contained. Then, $\mathcal{H}_{m}=\bigcup_{k=1}^{m}\mathcal{E}_{m,k}$. It can be easily verified that the size of the ZDD representing $Z(\mathcal{E}_{m,k})$ is $O(m^{2})$ for any order of elements (see Section 3.4). However, it is known that the ZDD representing $\mathcal{H}_{m}$ must become exponential in size.
Theorem 3.2 ([3]).
For any order $<$ of elements, $B_{<}(\mathcal{H}_{m})=\Omega(2^{m/5})$. Thus, by Lemma 2.2, $Z_{<}(\mathcal{H}_{m})=\Omega(2^{m/5}/m)$.
Definition 3.3.
@@
For $k=1,\ldots,m$, let $\mathcal{Q}_{m,k}$ be the family of subsets of $\{y_{1},\ldots,y_{m^{2}}\}$ such that there is exactly one element from $y_{m(k\!-\!1)\!+\!1},y_{m(k\!-\!1)\!+\!2},\ldots,y_{m(k\!-\!1)\!+\!m}$, and let $\mathcal{Q}_{m,m+k}$ be those such that there is exactly one element from $y_{k},y_{m\!+\!k},\ldots,y_{m(m\!-\!1)\!+\!k}$. Then, $\mathcal{P}_{m}=\bigcap_{k=1}^{2m}\mathcal{Q}_{m,k}$. Here, $Z(\mathcal{Q}_{m,k})=O(m^{2})$ for any order of elements, as proved in Section 3.4. However, it is again proved that the ZDD representing $\mathcal{P}_{m}$ must become exponential in size.
Theorem 3.4 ([13, Theorem K]).
For any order $<$ of elements, $B_{<}(\mathcal{P}_{m})=\Omega(m2^{m})$. Thus, by Lemma 2.2, $Z_{<}(\mathcal{P}_{m})=\Omega(2^{m}/m)$.
We first show the exponential blow-up cases for a specific order of elements in Section 3.2. However, we see that the size of ZDD representing the hidden weighted bit function or the permutation function is exponential regardless of the order of elements. Therefore, in Section 3.3, we prove that for each family generated by the operation in Section 3.2, the ZDD size representing it remains exponential regardless of the order of elements. This means that for each operation, there exists an instance in which the input ZDD size can be polynomial by managing the element order but the output ZDD size must be exponential for any order. Section 3.4 completes the proof by showing that some families can be represented by polynomial-sized ZDDs. Finally, Section 3.5 gives some discussions on the obtained result.
@@ 3.2.1 Join, Disjoint Join, Joint Join, Meet, and Delta
For these operations, we constitute a pair of families that incur the union of $O(m)$ subfamilies. Combined with $\mathcal{E}_{m,k}$, the result after taking an operation contains $\bigcup_{k}\mathcal{E}_{m,k}=\mathcal{H}_{m}$, which is the hidden weighted bit function for which the ZDD size is exponential in $m$.
Theorem 3.5.
@@
Figure 2: Example of blow-up for join (left) and quotient (right) operations. Blue triangles mean that the ZDD size representing this family is polynomial in $m$, while red triangle means that its size is exponential in $m$. Arcs going to $\bot$ terminal are omitted.
Proof 3.6.
Let us consider the families of subsets of $X\cup Y$, where $X\coloneqq\{x_{1},\ldots,x_{m}\}$ and $Y\coloneqq\{y_{1},\ldots,y_{m}\}$. We determine the order of elements as $x_{1},\ldots,x_{m},y_{1},\ldots,y_{m}$. We define $\mathcal{F}_{m}$ as
@@
$\mathcal{F}_{m}\sqcup\mathcal{G}_{m}$ $\textstyle=(\bigcup_{k=1}^{m}(\{\{x_{k}\}\}\sqcup\mathcal{E}_{m,k}))\sqcup\{X\}=\bigcup_{k=1}^{m}((\{\{x_{k}\}\}\sqcup\mathcal{E}_{m,k})\sqcup\{X\})$
$\textstyle=\bigcup_{k=1}^{m}(\{X\}\sqcup\mathcal{E}_{m,k})=\{X\}\sqcup(\bigcup_{k=1}^{m}\mathcal{E}_{m,k})=\{X\}\sqcup\mathcal{H}_{m},$
@@
$\mathcal{F}_{m}\operatorname{\dot{\bowtie}}\mathcal{G}_{m}$ $\textstyle=\bigcup_{k=1}^{m}((\{x_{k}\}\cup(X\setminus\{x_{k}\}))\sqcup\mathcal{E}_{m,k})=\{X\}\sqcup(\bigcup_{k=1}^{m}\mathcal{E}_{m,k})=\{X\}\sqcup\mathcal{H}_{m},$
@@
For the delta operation, we let $\mathcal{G}_{m}=2^{X}$. Since $\{\{x_{k}\}\}\operatorname{\boxplus}2^{X}=2^{X}$ for any $k$, we have
$\mathcal{F}_{m}\operatorname{\boxplus}\mathcal{G}_{m}$ $\textstyle=\bigcup_{k=1}^{m}((\{\{x_{k}\}\}\operatorname{\boxplus}2^{X})\sqcup\mathcal{E}_{m,k})=2^{X}\sqcup(\bigcup_{k=1}^{m}\mathcal{E}_{m,k})=2^{X}\sqcup\mathcal{H}_{m}.$
The ZDD size of $\mathcal{F}_{m}\operatorname{\boxplus}\mathcal{G}_{m}$ is at least $Z(\mathcal{H}_{m})=\Omega(2^{m/5}/m)$.
@@ 3.2.2 Quotient and Remainder
For the quotient operation, we constitute a pair of families such that performing an operation incurs the intersection of $O(m)$ subfamilies. Here, let $\mathcal{E}^{\prime}_{m,k}\coloneqq 2^{Y}\setminus\mathcal{E}_{m,k}$ be the complement of $\mathcal{E}_{m,k}$ regarding the family of subsets of $Y$. By De Morgan’s laws, we have $\bigcap_{k}\mathcal{E}^{\prime}_{m,k}=2^{Y}\setminus(\bigcup_{k}\mathcal{E}_{m,k})=2^{Y}\setminus\mathcal{H}_{m}\eqqcolon\mathcal{H}^{\prime}_{m}$. The ZDD size representing $\mathcal{H}^{\prime}_{m}$ can be lower bounded by the following lemma.
Lemma 3.7.
@@
Proof 3.8 (Proof of Lemma 3.7).
$\mathcal{F}\supseteq\mathcal{G}$ implies $\mathcal{F}\setminus(\mathcal{F}\setminus\mathcal{G})=\mathcal{G}$. Since the ZDD size after taking the difference can be bounded by the product of the sizes of operand ZDDs, we have $Z(\mathcal{G})=O(Z(\mathcal{F})Z(\mathcal{F}\setminus\mathcal{G}))$. Suppose $Z(\mathcal{F}\setminus\mathcal{G})=o(g(m)/f(m))$. Then, $Z(\mathcal{G})=o(f(m)\cdot(g(m)/f(m)))=o(g(m))$, refuting the assumption $Z(\mathcal{G})=\Omega(g(m))$. Therefore, $Z(\mathcal{F}\setminus\mathcal{G})=\Omega(g(m)/f(m))$.
Since $Z(2^{Y})=O(m)$ and $Z(\mathcal{H}_{m})=\Omega(2^{m/5}/m)$, we have $Z(\mathcal{H}^{\prime}_{m})=\Omega(2^{m/5}/m^{2})$.
Theorem 3.9.
@@
Proof 3.10.
@@
We have $Z(\mathcal{E}^{\prime}_{m,k})=O(m^{2})$ as proved in Section 3.4, and thus $Z(\mathcal{F}_{m})=O(m^{3})$. We also define $\mathcal{G}_{m}\coloneqq\{\{x_{1}\},\ldots,\{x_{m}\}\}$, where $Z(\mathcal{G}_{m})=O(m)$.
@@
For the remainder operation, we prepared the same families. Since $\mathcal{G}_{m}\sqcup(\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m})=\{\{x_{1}\},\ldots,\{x_{m}\}\}\sqcup\mathcal{H}^{\prime}_{m}$, $Z(\mathcal{G}_{m}\sqcup(\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m}))=\Omega(2^{m/5}/m^{2})$. Also, since $S\in\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m}$ if and only if $S\cup G\in\mathcal{F}_{m}$ for all $G\in\mathcal{G}_{m}$, all of the subsets in $\mathcal{G}_{m}\sqcup(\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m})$ are also contained in $\mathcal{F}_{m}$. In other words, $\mathcal{F}_{m}\supseteq\mathcal{G}_{m}\sqcup(\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m})$. Therefore, by using Lemma 3.7, $Z(\mathcal{F}_{m}\operatorname{\%}\mathcal{G}_{m})=\Omega((2^{m/5}/m^{2})/m^{3})=\Omega(2^{m/5}/m^{5})$.